# Benchmarking Retrieval-Augmented Large Language Models in Biomedical NLP: Application, Robustness, and Self-Awareness

### v2

@@ 3.2 Comparison between RALs with backbone LLMs
Table 3: Results of various approaches for link prediction, text classification, question answering, and natural language inference. Underline with green shade indicates the best performance on each dataset.
ADE ChemProt GIT PHarmKG Hetionet
Corpus Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Unlabeled corpus 9.76 9.53 9.65 75.56 69.14 72.21 2.29 0.65 1.01 97.20 97.20 97.20 78.60 78.60 78.60
Labeled corpus 36.07 36.06 36.06 87.13 86.68 86.91 75.65 74.84 75.24 98.00 98.00 98.00 90.40 90.40 90.40
 None 34.94 34.79 34.86 42.60 41.51 42.05 42.60 41.51 42.05 97.60 97.60 97.60 78.54 78.54 78.54
Table 4: RAL Performance of ADE, ChemProt, GIT, PHarmKG and Hetionet on Testbed1: Unlabeled Robustness. On Testbed1, the used RAL is the one demonstrating the best performance on each dataset. Such as, on ADE, RAL is LLaMA2-13B with Contriever. Labeled corpus is the training set for each dataset. The unlabeled corpus refers to a training set devoid of labels for each dataset. Green shade refers to the best performance.
Ade-corpus-v2 SemClass MedMCQA BioNLI
Corpus Precision Recall F1 Precision Recall F1 Precision Recall F1 F1
Unlabeled corpus 94.80 94.80 94.80 6.83 6.83 6.83 35.19 35.27 35.19 10.91
Labeled corpus 96.80 96.80 96.80 79.33 79.33 79.33 40.38 40.49 40.42 69.21
 None 96.40 96.40 96.40 77.66 77.66 77.66 41.63 41.52 41.52 62.62
Table 5: RAL Performance of Ade-corpus-v2, SemClass, MedMCQA, and BioNLI on Testbed1: Unlabeled Robustness.
3.3 Results and Discussion on Testbed1: Unlabeled Robustness
@@
We evaluate the model performance based on the unlabeled corpus, and the results are shown in Table 4 and Table 5. We have the following observations:
@@
3.3.1 Error Analysis
To better understand the impact of the unlabeled corpus on model generation, this section primarily analyzes the RAL performance on ADE, GIT, and BioNLI, which exhibited the poorest performance among the nine datasets used. We primarily summarize two error types as shown in Table 6. We observed that with the unlabeled corpus, RAL tends to generate redundant information and struggles to accurately predict the output, such as the head entity or relation type in the triple extraction task.
Error type Dataset Input sentence Expected output Error output
Redundant information ADE the fourth patient showed rls symptoms that were initially caused by a 20-mg daily olanzapine dosage and were later mitigated when olanzapine was reduced and ropinirole was administered. {olanzapine, dosage, 20-mg daily} {olanzapine, dosage, rls symptoms that were initially caused by a 20-mg dail}
GIT inactivation kinetics of vacterial glycerol dehydratase (ec 4.2.1.30) in the course of its reaction with adenosylcobalamin (adocbl) and its analogs were investigated.. ground tail entity: glycerol dehydratase generated tail entity: adenosylcobalamin.. retrieved sentence: glycerol dehydratase
BIONLI negative negative retrieved sentence..
Error generation ADE four patients receiving high-dose tamoxifen for greater than 1 year have demonstrated similar retinal changes. (tamoxifen, dosage, high-dose) (tamoxifen, effect, retinal changes.)
GIT inactivation of serum alkaline phosphatase by adrenaline and related substances . (adrenaline, inhibits, alkaline phosphatase) (alkaline phosphatase, interacts with, adrenaline)
BIONLI positive negative
3.4 Results and Discussion on Testbed2: Counterfactual Robustness
ADE ChemProt GIT PHarmKG Hetionet
Corpus Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Counterfactual corpus (20%) 41.41 41.31 41.36 74.22 66.90 70.37 74.50 71.61 73.03 97.40 97.40 97.40 94.80 94.80 94.80
Counterfactual corpus (80%) 41.94 36.90 39.26 82.96 33.57 47.79 74.84 74.84 74.84 97.80 97.80 97.80 85.26 85.26 85.26
Counterfactual corpus (100%) 25.60 25.59 25.59 84.12 83.07 83.59 75.22 74.41 74.81 97.60 97.60 97.60 76.60 76.60 76.60
Labeled corpus 36.07 36.06 36.06 87.13 86.68 86.91 75.65 74.84 75.24 98.00 98.00 98.00 90.40 90.40 90.40
 None 34.94 34.79 34.86 42.60 41.51 42.05 42.60 41.51 42.05 97.60 97.60 97.60 78.60 78.60 78.60
Table 7: RAL Performance of ADE, ChemProt, GIT, PHarmKG and Hetionet on Testbed2: counterfactual robustness. On Testbed2, the used RAL is the one demonstrating the best performance on each dataset. Such as, on ADE, RAL is LLaMA2-13B with Contriever. Labeled corpus is the training set for each dataset. The 20/80/100% denotes a labeled corpus where 20/80/100% of instances are mislabeled for each dataset. Green shade refers to the best model performance.
Ade-corpus-v2 SemClass MedMCQA BioNLI
Corpus Precision Recall F1 Precision Recall F1 Precision Recall F1 F1
Counterfactual corpus (20%) 95.80 95.80 95.80 73.33 73.33 73.33 34.93 35.03 34.94 47.76
Counterfactual corpus (80%) 95.00 95.00 95.00 75.66 75.66 75.66 36.67 36.44 36.47 64.63
Counterfactual corpus (100%) 96.80 96.80 96.80 77.66 77.66 77.66 37.27 37.40 37.28 46.46
Labeled corpus 96.80 96.80 96.80 79.33 79.33 79.33 40.38 40.49 40.42 69.21
 None 96.40 96.40 96.40 77.66 77.66 77.66 35.19 35.27 35.19 62.62
Table 8: RAL Performance of Ade-corupus-v2, semCLass, MedMCQA, and BIoNL on Testbed2 counterfactual robustness.
@@ 3.4 Results and Discussion on Testbed2: Counterfactual Robustness
We evaluate the model performance based on different counterfactual rates, and the results are shown in Table 7 and Table 8. We have the following observations:
@@
(1) Counterfactual corpus posses a challenge for RALs. On ChemProt, counterfactual instances significantly influence the model performance. For instance, when the counterfactual rate is set to 80%, the triple F1 drops to 47.79%, showcasing a considerable disparity compared to the triple F1 performance on the labeled corpus. Similar observations are noted in GIT, PharmKG, ADE-corpus-v2, SemClass, and BioNLI. This suggests that RALs can be easily misled by counterfactual corpus.
@@
(2) A lower counterfactual rate may have a reduced impact on RALs. On ADE and Hetionet, we observed that when the counterfactual corpus is set to 20%, the model performance is better than the factual corpus. We speculate that retrievers have a greater chance of obtaining useful information when the counterfactual rate is lower.
@@
(3) The counterfactual corpus can still contribute to improving LLM performance. On ADE, ChemProt, GIT, PHarmKG, Hetionet, Ade-corpus-v2, SemClass, MEdMCQA, BIoNLI. The interesting finding is that even with a counterfactual corpus, the RAL performance often surpasses the original LLM. We speculate that the counterfactual corpus may have a beneficial effect on LLMs. Despite the content of the instances being counterfactual, the provided templates still aid in generation.
@@
3.5 Results and Discussion on Testbed3: Diverse Robustness
ADE ChemProt GIT PHarmKG Hetionet
Corpus Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Diverse corpus 11.21 9.15 10.08 78.56 76.41 77.47 74.88 65.38 69.80 97.20 97.20 97.20 75.41 75.41 75.41
Labeled corpus 36.07 36.06 36.06 87.13 86.68 86.91 75.65 74.84 75.24 98.00 98.00 98.00 90.40 90.40 90.40
 None 34.94 34.79 34.86 42.60 41.51 42.05 42.60 41.51 42.05 97.60 97.60 97.60 78.60 78.60 78.60
Table 9: RAL Performance of ADE, ChemProt, GIT, PHarmKG and Hetionet on Testbed3: diverse robusteness. On Testbed3, the used RAL is the one demonstrating the best performance on each dataset. Such as, on ADE, RAG is LLaMA2-13B with Contriever. Labeled corpus is the training set for each dataset. The diversity labeled corpus of n refers to a collection of (0,..,n-1) corpus.
Ade-corpus-v2 SemClass MedMCQA BioNLI
Corpus Precision Recall F1 Precision Recall F1 Precision Recall F1 F1
Diversity labeled corpus 96.20 96.20 96.20 75.33 75.33 75.33 25.01 25.02 24.87 80.32
Labeled corpus 96.80 96.80 96.80 79.33 79.33 79.33 40.38 40.49 40.42 69.21
 None 96.40 96.40 96.40 77.66 77.66 77.66 41.63 41.52 41.52 62.62
Table 10: RAL Performance of Ade-corupus-v2, SemClass, MedMCQA, and BioNLI on Testbed 3: diverse robusteness.
We evaluate the model performance of diversity robustness, and the results are shown in Table 9 and Table 10. We have the following observations:
The diversity labeled corpus poses a challenge to improve RALs. We found that RALs consider the knowledge in the diverse corpus as noise, which could potentially impact RAL performance, particularly evident in ADE and MedMCQA datasets. However, on BioNLI, the diversity labeled corpus could contribute to enhancing the model performance. We speculate that one reason is the retriever we used couldn’t retrieve useful information, while another reason could be that the corpus lacks the necessary information.
3.5.1 Error Analysis
On ADE, we discovered that the Diversity-labeled corpus also leads to redundancy in RAL generation, for instance, in sentence easily reversible hypoxemia and hypotension induced by nimodipine., the expected tail entity is hypotension, while RAL regarded the hypoxemia and hypotension induced by nimodipine. as the entity. It also struggles with extracting complex entities. For example, in the sentence clinical, spectroscopic, and imaging abnormalities resolved with discontinuation of metronidazole, clinical, spectroscopic, and imaging abnormalities is considered the ground truth, while RAL regards the entire sentence clinical, spectroscopic, and imaging abnormalities resolved with discontinuation of metronidazole as a single entity. In summary, we find that the primary challenge lies in entity recognition, especially in the recognition of tail entities. On MedMCQA, we observed that error generation primarily stemmed from misjudgment. For instance, in sentence Question: All of the following muscles are elevators of the mandible EXCEPT: Options: (A) Digastric; (B) Masseter; (C) Medial pterygoid; (D) Temporalis, the ground truth is A, while RAL generates the D.
3.6 Results and Discussion on Testbed4: Negative Awareness

### v3

@@ 2.1 Results of RALs and backbone LLMs
Table 1: Micro F1 results for triple extraction, link prediction, and text classification, along with Macro-average F1 for question answering and BioNLI, across various approaches on 11 datasets.Underline indicates the best performance on each dataset. For complete results, please refer to the Supplementary Materials.
2.2 Results of Testbed 1, 2 and 3
Triple Extraction Link Prediction Classification NL Inference
Corpus ADE GIT PHarmKG Hetionet Ade-corpus-v2 SemClass SDoH classification BioNLI
Unlabeled corpus 14.39 1.01 97.20 78.60 93.00 6.83 72.99 61.38
Counterfactual corpus (20%) 45.06 73.03 97.40 94.80 95.80 73.33 75.50 87.62
Counterfactual corpus (80%) 30.07 74.84 97.80 85.26 95.00 75.66 13.34 60.09
Counterfactual corpus (100%) 39.98 74.81 97.60 76.60 96.80 77.66 0 58.52
Diverse corpus 10.08 69.80 97.20 75.41 96.20 75.33 83.90 80.32
Labeled corpus 45.65 75.24 98.00 90.40 96.80 79.33 81.43 91.19
None 26.76 42.05 97.60 78.54 96.40 77.66 74.68 63.20
Table 2: Micro F1 results for triple extraction, link prediction, text classification, along with Macro-average F1 for BioNLI, across various approaches on 8 datasets under Testbed 1, Testbed 2, and Testbed 3. Underline indicates the best performance on each dataset. Note: we utilize the best RAL in the table 1 as the backbone.
@@ 2.2 Results of Testbed 1, 2 and 3
We evaluate the model performance based on the unlabeled corpus, and the results are shown in Table 2. We have the following observations:
@@
We evaluate the model performance based on different counterfactual rates, and the results are shown in Table 2. We have the following observations:
@@
(1) Counterfactual corpus posses a challenge for RALs. On ADE, counterfactual instances significantly influence the model performance. For instance, when the counterfactual rate is set to 80%, the triple F1 drops to around 10%, showcasing a considerable disparity compared to the triple F1 performance on the labeled corpus. Similar observations are noted in GIT, PharmKG, ADE-corpus-v2, SemClass, and BioNLI. This suggests that RALs can be easily misled by counterfactual corpus.
@@
(2) A lower counterfactual rate may have a reduced impact on RALs. On Hetionet, we observed that when the counterfactual corpus is set to 20%, the model performance is better than the factual corpus. We speculate that retrievers have a greater chance of obtaining useful information when the counterfactual rate is lower.
@@
(3) The counterfactual corpus can still contribute to improving LLM performance. On ADE, GIT, PHarmKG, Hetionet, Ade-corpus-v2, SemClass, BIoNLI. The interesting finding is that even with a counterfactual corpus, the RAL performance often surpasses the original LLM. We speculate that the counterfactual corpus may have a beneficial effect on LLMs. Despite the content of the instances being counterfactual, the provided templates still aid in generation.
@@
We evaluate the model performance of diversity robustness, and the results are shown in Table 2. We have the following observations:
The diversity labeled corpus poses a challenge to improve RALs. We found that RALs consider the knowledge in the diverse corpus as noise, which could potentially impact RAL performance, particularly evident in ADE datasets. However, on BioNLI, the diversity labeled corpus could contribute to enhancing the model performance. We speculate that one reason is the retriever we used couldn’t retrieve useful information, while another reason could be that the corpus lacks the necessary information.
In addition to the public dataset, we also used our labeled private dataset, SDoH Classification, which is derived from real-world data. We observed the same conclusion as with the public dataset: the unlabeled dataset and counterfactual corpus influence the performance of RALs. We did not include the our labeled private dataset DS, as RALs without retrievers achieved better performance.
2.3 Results of Testbed4: Negative Awareness